Bert is 4 times as old as Harry. Three years ago he was 7 times as old as Harry. How old is each now?
Answer by jorel1380(3719)

Let h and 4h be Harry and Bert, respectively. Then:
4h-3=7(h-3)
4h-3=7h-21
3h=18
h=6
Harry is 6; Bert is 24
☺☺☺☺